# The Exact Uncertainty Relation and Geometric Speed Limits in Krylov Space

### v1

@@ Abstract
This work is dedicated to Sudipta Mukherji on the occasion of his 60th birthday.
@@ I Introduction
The uncertainty principle is a defining feature of quantum mechanics, limiting the precision with which certain pairs of observables can be simultaneously measured. Beyond the canonical position–momentum relation, energy–time uncertainty relations Mandelstam and Tamm (1945); Margolus and Levitin (1998) imply fundamental limits on how rapidly a quantum state may evolve, known as quantum speed limits (QSLs). Such limits play a central role in quantum control, quantum information processing, and quantum computation.
Most uncertainty relations take the form of inequalities and therefore yield lower bounds on the time required for a given evolution. In contrast, Hall derived a state-dependent exact uncertainty relation Hall (2001) for two observables $A$ and $B$ in a pure state $\rho=|\psi\rangle\langle\psi|$,
@@
where $\Delta B_{nc}$ denotes the nonclassical component of $B$, and $\delta_{B}A$ is a Fisher-information-like functional of $A$. Although explicitly state dependent, this relation yields an equality-type QSL once the nonclassical generator is identified in a suitable basis. Indeed, Ref. Pati et al. (2023) used Hall’s relation to construct an exact QSL for quantum dynamics.
In this work, we show that operator dynamics admits an exact geometric formulation in Krylov space, in which quantum speed limits arise as kinematic identities. Specifically, we demonstrate that Hall’s exact uncertainty relation acquires a particularly simple and universal form in the Krylov basis, where the nonclassical part of the Liouvillian coincides with the full generator. As an immediate consequence, Lanczos representation of the Liouvillian Lanczos (1950); Viswanath and Müller (2008) reveals operator evolution on the unit Krylov sphere with a constant speed fixed solely by the first Lanczos coefficient. This could be identified as a geometric Lieb–Robinson–type bound intrinsic to Krylov dynamics which remains exact even when the Lanczos coefficients grow without bound, as in typical chaotic systems. In our work, we demonstrate this by introducing the motion of amplitude fronts in Krylov index space which is distinct from the geometric motion of the operator state on the Krylov sphere. We show that even when the operator front accelerate rapidly in Krylov index space, the actual geometric motion of the operator state remains strictly bounded, thereby resolving the apparent tension between fast Krylov growth and finite causal structure.
@@ III.1 Geometry of Krylov Sphere
where we have also used that fact that $\dot{\Phi}\cdot\ddot{\Phi}=0$ and $|\ddot{\Phi}|^{2}=b_{1}^{2}(b_{1}^{2}+b_{2}^{2})$. One may also define the torsion using the triple of derivatives  [^2] Practically, the determinant can be computed by forming the $3\times 3$ Gram matrix $G_{ij}=v_{i}\!\cdot\!v_{j}$ with $v_{1}=\dot{\Phi},v_{2}=\ddot{\Phi},v_{3}=\dddot{\Phi}$ and then using $\det[\dot{\phi},\ddot{\phi},\dddot{\phi}]^{2}=\det G$.
@@ III.2 Light cone in the Krylov sphere
The resulting wavefunction is peaked around $n\approx 2bt$ and decays exponentially outside this region, demonstrating that the Krylov light–cone bound in index space is essentially sharp.
@@
The resolution of this apparent paradox is conceptual rather than technical. It arises from distinguishing two fundamentally different notions of motion: (i) the geometric motion of the operator state on the Krylov sphere, and (ii) the motion of the amplitude “front” in Krylov index space  [^5] The notion of a propagating wavefront in Krylov index space and its accelerated motion for growing Lanczos coefficients has appeared previously in studies of Krylov complexity and operator growth, see e.g. Refs. Rabinovici et al. (2023); Ambrosini et al. (2025). Our emphasis here is different: we distinguish this index-space redistribution from the true geometric motion of the operator state, which we show is exactly bounded by a linear arc-length law controlled solely by the first Lanczos coefficient. . The measure of (i) is the true physical displacement of the operator along a unit-speed trajectory on the Krylov sphere and is governed by the arc–length $\ell_{K}(t)$. This motion therefore defines a natural Lieb–Robinson–type velocity in Krylov space, $v_{\mathrm{op}}=b_{1}$. By contrast, (ii) merely tracks how the wavefunction is distributed among basis elements labeled by the discrete index $n$. Clearly, these two notions need not scale in the same way. In particular, when the Lanczos coefficients grow with $n$, successive Krylov levels correspond to progressively smaller angular displacements on the sphere. As a result, the operator may traverse many Krylov levels while covering only a small geometric distance. Thus, even when the front $n(t)$ accelerates, the geometric motion remains strictly bounded by $\ell_{K}(t)=b_{1}t$, and no causal constraint is violated.
In order to demonstrate this, let us note once again that the Lanczos recursion couples only nearest neighbors, and therefore the characteristic time to transfer amplitude from level $n$ to $n+1$ can be estimated as $\Delta t_{n}\sim 1/{b_{n}}$. This immediately implies the minimal-time constraint for reaching level n
@@
For bounded $b_{n}$ this yields linear growth $n(t)\leq b_{1}t$, while unbounded $b_{n}$ leads to nonlinear fronts, e.g. (Further details are given in Supplementary materials),
@@
Thus, even when Lanczos coefficients diverge and the front accelerates, the operator’s geometric displacement remains strictly bounded by $\ell_{K}(t)=b_{1}t$. The true light cone is linear and controlled solely by $b_{1}$, resolving the apparent conflict between unbounded Lanczos growth and finite causal structure.
@@ Acknowledgements
We would like to thank Mohammad Reza Tanhayi, Mohammad Javad Vasli for many insightful discussions on various aspects of complexity. We would like to thank Arun K. Pati for his comments on an earlier version of this manuscript. The work of M. A. is supported by the Iran National Science Foundation (INSF) under Project No. 4023620.
@@ Appendix A Hall uncertainty
It is worth noting that the price we paid to get an equality is that the uncertainty relation (1) is state dependent which may not be so restrictive, as when it comes to an actual experiment we always start with a prepared quantum state. Moreover, for non-pure states the above uncertainty relation also reduces to an inequality Hall (2001).
@@
vanishes identically. Thus ${\cal L}_{nc}={\cal L}$ and using (7)
@@ Appendix B Geodesic and Non-geodesic Operator Trajectories on
and initial operator ${\cal O}=\sigma_{x}$. One finds that the Krylov subspace is two-dimensional, spanned by $\{\sigma_{x},\sigma_{y}\}$. The Lanczos coefficients are $b_{1}=\omega$ and $b_{n}=0$ for $n\geq 2$, and the Schrödinger equations become
@@
With $\phi_{0}^{2}+\phi_{1}^{2}=1$, this describes uniform rotation on the unit circle with angular velocity $\omega$. Form this simple example emerges a geometric picture for operator evolution in which the “operator particle” moves along a great circle in the Krylov space with constant speed $b_{1}=\omega$.
In this example since only $b_{1}$ was non-zero the operator moves along the geodesic. To see how non-geodesic motion occurs let us modified the above Hamiltonian by adding a transverse magnetic filed along the $x$ direction
@@
with seed operator $\mathcal{O}(0)=\sigma_{x}$. The orthonormal Krylov basis are

### v2

@@ I Introduction
The uncertainty principle is a defining feature of quantum mechanics, limiting the precision with which certain pairs of observables can be simultaneously measured. Beyond the canonical position–momentum relation, the energy–time uncertainty relations Mandelstam and Tamm (1945); Margolus and Levitin (1998) imply fundamental limits on how rapidly a quantum state may evolve, known as quantum speed limits (QSLs). Such limits play a central role in quantum control, quantum information processing, and quantum computation.
Most uncertainty relations take the form of inequalities, and therefore yield lower bounds on the time required for a given evolution. In contrast, Hall derived a state-dependent exact uncertainty relation Hall (2001) for two observables $A$ and $B$ in a pure state $\rho=|\psi\rangle\langle\psi|$,
@@
where $\Delta B_{nc}$ denotes the nonclassical component of $B$, and $\delta_{B}A$ is a Fisher-information-like functional of $A$. Although explicitly state dependent, this relation yields an equality-type QSL once the nonclassical generator is identified in a suitable basis. Indeed, in Pati et al. (2025); Srivastav et al. (2025) Hall’s relation was used to construct an exact QSL for quantum dynamics.
In this work, we show that operator dynamics admits an exact geometric formulation in Krylov space, in which quantum speed limits arise as kinematic identities. Specifically, we demonstrate that Hall’s exact uncertainty relation acquires a particularly simple and universal form in the Krylov basis, where the nonclassical part of the Liouvillian coincides with the full generator. As an immediate consequence, Lanczos representation of the Liouvillian Lanczos (1950); Viswanath and Müller (2008) reveals operator evolution on the unit Krylov sphere with a constant speed fixed solely by the first Lanczos coefficient. This can be identified as a geometric Lieb–Robinson–type bound intrinsic to Krylov dynamics which remains exact even when the Lanczos coefficients grow without bound, as in typical chaotic systems. In our work, we demonstrate this by introducing the motion of amplitude fronts in Krylov index space, which is distinct from the geometric motion of the operator state on the Krylov sphere. We show that even when the operator front accelerate rapidly in Krylov index space, the actual geometric motion of the operator state remains strictly bounded, thereby resolving the apparent tension between fast Krylov growth and finite causal structure.
@@ III.1 Geometry of Krylov Sphere
where we have also used the fact that $\dot{\Phi}\cdot\ddot{\Phi}=0$ and $|\ddot{\Phi}|^{2}=b_{1}^{2}(b_{1}^{2}+b_{2}^{2})$. One may also define the torsion using the triple of derivatives  [^2] Practically, the determinant can be computed by forming the $3\times 3$ Gram matrix $G_{ij}=v_{i}\!\cdot\!v_{j}$ with $v_{1}=\dot{\Phi},v_{2}=\ddot{\Phi},v_{3}=\dddot{\Phi}$ and then using $\det[\dot{\phi},\ddot{\phi},\dddot{\phi}]^{2}=\det G$.
@@ III.2 Light cone in the Krylov sphere
The resulting wavefunction is peaked around $n\approx 2bt$ and decays exponentially outside this region, demonstrating that the Krylov light–cone bound in the index space is essentially sharp.
@@
The resolution of this apparent paradox is conceptual rather than technical. It arises from distinguishing two fundamentally different notions of motion: (i) the geometric motion of the operator state on the Krylov sphere, and (ii) the motion of the amplitude “front” in Krylov index space  [^5] The notion of a propagating wavefront in Krylov index space and its accelerated motion for growing Lanczos coefficients has appeared previously in studies of Krylov complexity and operator growth, see e.g. Refs. Rabinovici et al. (2023); Ambrosini et al. (2025); Hörnedal et al. (2023); Nandy et al. (2025). Our emphasis here is different: we distinguish this index-space redistribution from the true geometric motion of the operator state, which we show is exactly bounded by a linear arc-length law controlled solely by the first Lanczos coefficient. . The measure of (i) is the true physical displacement of the operator along a unit-speed trajectory on the Krylov sphere and is governed by the arc–length $\ell_{K}(t)$. This motion therefore defines a natural Lieb–Robinson–type velocity in Krylov space, $v_{\mathrm{op}}=b_{1}$. By contrast, (ii) merely tracks how the wavefunction is distributed among basis elements labeled by the discrete index $n$. Clearly, these two notions need not scale in the same way. In particular, when the Lanczos coefficients grow with $n$, successive Krylov levels correspond to progressively smaller angular displacements on the sphere. As a result, the operator may traverse many Krylov levels while covering only a small geometric distance. Thus, even when the front $n(t)$ accelerates, the geometric motion remains strictly bounded by $\ell_{K}(t)=b_{1}t$, and no causal constraint is violated.
In order to demonstrate this, let us note once again that the Lanczos recursion couples only nearest neighbors, and therefore the characteristic time to transfer the amplitude from level $n$ to level $n+1$ can be estimated as $\Delta t_{n}\sim 1/{b_{n}}$. This immediately implies the minimal-time constraint for reaching level n
@@
For bounded $b_{n}$ this yields linear growth $n(t)\leq b_{1}t$, while unbounded $b_{n}$ leads to nonlinear fronts, e.g. (Further details are given in the supplemental materials),
@@
Thus, even when the Lanczos coefficients diverge and the front accelerates, the geometric displacement of the operator remains strictly bounded by $\ell_{K}(t)=b_{1}t$. The true light cone is linear and is controlled solely by $b_{1}$, resolving the apparent conflict between unbounded Lanczos growth and a finite causal structure.
@@ Acknowledgements
We would like to thank Mohammad Reza Tanhayi, Mohammad Javad Vasli for many insightful discussions on various aspects of complexity. We would like to thank Arun K. Pati for his comments on an earlier version of this manuscript. The work of M. A. is supported by the Iran National Science Foundation (INSF) under Project No. 4023620. This work is dedicated to Sudipta Mukherji on the occasion of his 60th birthday.
@@ Appendix A Hall uncertainty
It is worth noting that the price we paid to get an equality is that the uncertainty relation (S1) is state dependent which may not be so restrictive, as when it comes to an actual experiment we always start with a prepared quantum state. Moreover, for non-pure states the above uncertainty relation also reduces to an inequality Hall (2001).
@@
vanishes identically. Thus ${\cal L}_{nc}={\cal L}$ and using equation (7)
@@ Appendix B Geodesic and Non-geodesic Operator Trajectories on
and initial operator ${\cal O}=\sigma_{x}$. The Krylov subspace turns out to be two-dimensional, spanned by $\{\sigma_{x},\sigma_{y}\}$. The Lanczos coefficients are $b_{1}=\omega$ and $b_{n}=0$ for $n\geq 2$, and the Schrödinger equations become
@@
With $\phi_{0}^{2}+\phi_{1}^{2}=1$, this describes uniform rotation on the unit circle with angular velocity $\omega$. From this simple example emerges a geometric picture for operator evolution in which the “operator particle” moves along a great circle in the Krylov space with constant speed $b_{1}=\omega$.
In this example since only $b_{1}$ was non-zero the operator moves along the geodesic. To see how non-geodesic motion occurs, let us modify the above Hamiltonian by adding a transverse magnetic field along the $x$ direction
@@
with seed operator $\mathcal{O}(0)=\sigma_{x}$. The orthonormal Krylov basis is